Suppose that a firm's labor demand curve is given by w = 20 - L. Currently, a firm pays its workers $10. The production function for the firm is y = 4L, where y is the output. The firm sells its product for $3.

The firm decides to shift to an efficiency wage model, where they pay workers $15. The new production function with the efficiency wage is y = 2L^2. The firm still sells its product for $3. How much higher are profits for the firm with the efficiency wage, compared to without?


1
Expert's answer
2021-04-19T07:43:42-0400

If w = 20 - L, and w = $10, then:

10 = 20 - L,

L = 10 units.

y = 4×10 = 40 units.

Total profit was:

TP = 3×40 - 10×10 = $20.

If new wage is w = $15, then:

15 = 20 - L,

L = 5 units.

"y = 2\u00d75^2 = 50" units.

TP = 3×50 - 5×15 = $75.

So, the profit will increase by 75 - 20 = $55.